Determine the pH of the following solution:

0.852 M LiC2H3O2
Answer #1

C2H3O2- (acetate ion) partly hydrolysed by water

C2H3O2-(aq) + H2O(l) <--------> HC2H3O2(aq) + OH-(aq)

Kb = [HC2H3O2][OH​​​​​​-] / [C2H3O2-] = 5.56 ×10-10

Initial concentration

[C2H3O2- ] = 0.852

[HC2H3O2] = x

[ OH-] = x

change in concentration

[C2H3O2-] = -x

[HC2H3O2] = +x

[OH-] = + x

equilibrium concentration

[C2H3O2-] = 0.852 - x

[HC2H3O2] = x

[OH-] = x

so,

x2/( 0.852 - x) = 5.56 ×10-10

we can assume 0.852 - x = 0.852 because x is small value

x2/ 0.852 = 5.56 ×10-10

x2 = 4.737 × 10-10

  x = 2.176 × 10-5

[OH-] = 2.176 ×10-5M

pOH = -log[OH-]

pOH = - log( 2.176 ×10-5)

pOH = 4.66

pH = 14 - 4.66

pH = 9.34